# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

### v1

@@ 5.2 Locally constant Lagrange multiplier
In this subsection, we essentially follow McCann’s arguments [31, Section 5] but with some refinements to prove “weak” versions of Theorem 2.25 Parts (iviv-$v$). The “weak” version serves as the bridge linking the norm of $\rho$ and the norm of potential $V_{\rho}$, which can help to show the global continuity of the $W^{\infty}$ local minimizer $\rho$ as we have seen in the proof of Theorem 2.25
@@
We refer to [31], but slightly modify the arguments and provide a more detailed account (see Remark 5.9). Define the convex cone $P_{loc}:=\left\{\sigma\in P_{\infty}\mid\text{spt }\sigma\subset M\right\}$, which is a cone, and let $U=\{\rho\in X\mid\rho\geq 0,U(\rho)<\left.\infty,G(\rho,\rho)<\infty,T_{J}(\rho)<\infty\right\}$ be the same as in subsection 5.1. Thanks to Lemma 5.3 and Lemma 5.7, we have $\rho\in U,E_{J}(\rho)$ is $P_{loc}$-differentiable at $\rho$. Given $\kappa\in W_{loc}={R}\left(\mathbb{R}^{3}\right)\cap\{\rho+\sigma\in U\mid\sigma\in P_{loc}\}$, where the mass constraint is satisfied, thanks to Lemma 4.7 ($i$) we have $W^{\infty}(\rho,\kappa)\leq\operatorname{diam}M\leq 2\delta$, which shows that $\rho$ minimizes $E_{J}(\kappa)$ on $W_{loc}$. In particular, $\rho$ is a weak local minimizer restricted on $W_{loc}$, which is defined in Appendix A.1. Moreover, since the open set $M$ intersects spt $\rho$, it must carry positive mass under $\rho$, see for example [17, Section 3.4, Exercise 25]. Therefore, similar as Remark 4.11, there is a smaller subset $C\subset M$ of positive measure on which $\rho(x)$ is bounded away from zero and infinity. Let $\mathbf{1}_{C}$ is the indicator function of set $C$, then by construction we have both $\pm\mathbf{1}_{C}\in P_{loc}$. For all $\sigma\in P_{loc}$, let $w=\sigma-\frac{\int\sigma}{\int\mathbf{1}_{C}}\mathbf{1}_{C}$, thanks to Remark 4.12, we know $\rho+tw\in W_{loc}$ when $t$ is small enough. Thanks to Theorem A.4 and Remark A.5, the conditions above imply that there is a unique Lagrange multiplier $\lambda\in R$ such that $\forall\sigma\in P_{loc}$, we have
$\int E_{J}^{\prime}(\rho)\sigma\geq\lambda\int\sigma$ (31)
Claim: (1) $E_{J}^{\prime}(\rho)(x)=A^{\prime}(\rho(x))-V_{\rho}(x)-\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))\geq\lambda$ on $M$ a.e. and (2) $E_{J}^{\prime}(\rho)=\lambda$ on $M\cap\{x\mid\rho(x)>0\}$ a.e.
@@
In fact, If $E_{J}^{\prime}(\rho)<\lambda$ on a subset $K\subset M$ which had positive measure, similar as Remark 4.10, this subset may be taken slightly smaller so that $\rho$ is bounded on $K$. Then we take $\sigma=\mathbf{1}_{K}\in P_{loc}$ and would see it contradict (31), thus Claim (1) holds true. On the other hand, if $E_{J}^{\prime}(\rho)>\lambda$ on a subset $K\subset M$ with positive measure and $\rho>0$ in $K$, then similarly $K$ may be taken slightly smaller so that $\rho$ is bounded away from zero and infinity on $K$; in this case $\mathbf{1}_{K}\in P_{loc}$ can make a contradiction to (31) again, thus Claim (2) holds true.
@@
By Remark 2.5 we know $A^{\prime}(0)=0$, and $A^{\prime}(s)>0$ when $s>0$. If $V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda>0$, then by Claim (1) $A^{\prime}(\rho(x))\geq V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda>0$, which implies $\rho>0$. By Claim (2) we can replace inequality by equation and get $A^{\prime}(\rho(x))=V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda$. If $V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\leq 0$, then $\rho=0$, otherwise $\rho>0$ implies again $0<A^{\prime}(\rho(x))=V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\leq 0$, which leads to a contradiction. Therefore, we have (EL), that is $A^{\prime}(\rho(x))=\left[V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\right]_{+}$ holds for almost all $x\in M$. Since $M$ has positive measure and $\rho$ in $M$ has positive mass, we know there exists a $x_{0}$ such that $\rho(x_{0})>0$ and thus $A^{\prime}(\rho(x_{0}))>0$, and (EL) holds at $x_{0}$. In particular, $0<A^{\prime}(\rho(x_{0}))=V_{\rho}(x_{0})+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x_{0}-\bar{x}(\rho))+\lambda$. It is clear that $\lambda$ is unique and given by this expression . ∎
@@
Compared to the McCann’s proof, we have added the following discussions:
@@
We recall again $\rho+tw\in{R}\left(\mathbb{R}^{3}\right)$ may not be $W^{\infty}$-continuous as function of $t$. In another word, whether a local minimizer is also a weak local minimizer remains unclear in general. However, at least we know a minimizer restricted to $W_{loc}$ is a weak local minimizer restricted to $W_{loc}$, so that we can apply Theorem A.4 and Remark A.5. This is why we first consider the local case Proposition 5.8 (the diameter of $M$ is small) before addressing the next case Proposition 5.11.
@@
Given $y\in\Omega_{i}$, we know the ball $B_{\delta}(y)$ intersects spt $\rho$. Thus Proposition 5.8 guarantees a unique $\lambda(y)$ such that (EL) holds a.e. on $B_{\delta}(y)$. The claim is that $\lambda(y)$ is independent of $y$. In fact, fix $y_{0}\in\Omega_{i}$, since $B_{\delta}(y_{0})$ is open, it will also be true that a slightly smaller ball $B_{\delta-\epsilon}(y_{0})$ intersects spt $\rho$. If $|x-y_{0}|<\epsilon$, then $M=B_{\delta}(x)\cap B_{\delta}(y_{0})$ intersects spt $\rho$ since $B_{\delta-\epsilon}(y_{0})\subset M$, and then we know $x\in\Omega_{i}$. $M$ has positive measure, thanks to Proposition 5.8, the uniqueness of $\lambda$ corresponding to $M$ forces $\lambda(x)=\lambda(y_{0})$. Thus $\lambda(y)$ is locally constant. As a result, the disjoint sets $C=\left\{x\in\Omega_{i}\mid\lambda(x)=\lambda(y_{0})\right\}$ and $D=\left\{x\in\Omega_{i}\mid\lambda(x)\neq\lambda(y_{0})\right\}$ are both open. Since $\Omega_{i}=C\cup D$ is connected and $y_{0}\in C$ implies $C$ is not empty, $C=\Omega_{i}$. Defining $\lambda_{i}:=\lambda(y_{0})$,(EL) must be satisfied a.e. on $\Omega_{i}$.
@@
Notice we discuss above the connected components of the $\delta$-neighbourhood of spt $\rho$. But the result can be easily converted to the case about $\delta$-neighbourhood of connected components of spt $\rho$.
@@
From the above proof, we can see the power of Wasserstein metric, which can extend the Euler-Lagrange equation (EL) to a $\delta$-neighborhood of connected component of spt $\rho$, which help us to show the global continuity of $\rho$ and then (EP’) holds as we discuss in Theorem 2.25. Moreover, it plays a role in the process of proving the Lagrange multiplier is negative. Actually, there is another way to show Lagrange multipliers should be negative, given by Auchmuty and Beals [5], with different condition.
@@ 5.3 Minimizers’ existence and nonexistence
We recall again under the topology induced by $W^{\infty}$ metric, it is not entirely certain whether a local minimizer is a weak local minimizer, see Remark 5.10. For comparison, we prove a local minimizer is a weak local minimizer under the topology inherited from topological vector space. Furthermore, we demonstrate that, in this case, a local minimum either does not exist or is not physically meaningful.
Denote $X=L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, $\widetilde{U}=\{\rho\in X\mid\rho\geq 0\}$. We first show in the topology inherited from the topological vector space, a local minimizer is a weak local minmizer.
@@
If the topology on ${R}\left(\mathbb{R}^{3}\right)$ is inherited from a topological vector space, and $\rho$ is a local minimizer for $E_{J}(\rho)$ restricted on ${R}\left(\mathbb{R}^{3}\right)$, then $\rho$ is a weak local minimizer restricted on ${R}\left(\mathbb{R}^{3}\right)$, which is defined in Appendix A.1.
@@
We denote $\left.N\left(\mathbb{R}^{3}\right)=\left\{{\sigma\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)}\right|\int\sigma=0\right\}$. We recall the setup in Appendix A.1, that $P(\rho)=\left\{{\sigma\in X}\right|\exists\epsilon(\sigma),~\text{such that}~\rho+t\sigma\in\widetilde{U},~0\leq t<\epsilon(\sigma)\}$.
Since the topology on $R(\mathbb{R}^{3})$ is inherited from a topological vector space, we know there is a topological vector space $Y$ with topology $\mathcal{T}$, such that $R(\mathbb{R}^{3})$ is a subset of $Y$, with the collection $\mathcal{T}_{R}=\left\{R\left(\mathbb{R}^{3}\right)\cap U\middle|U\in\mathcal{T}\right\}$ being the topology on $R(\mathbb{R}^{3})$. (See [33, Chapter 1, Section 16]).
Since $\rho$ is a local minimizer restricted on $R(\mathbb{R}^{3})$, then $\exists U_{R}\in\mathcal{T}_{R}$, such that $\rho\in U_{R}$ and $E_{J}(\rho)\leq E_{J}(\kappa)$ for all $\kappa\in U_{R}$. Notice that $U_{R}=R(\mathbb{R}^{3})\cap U$ for some $U\in\mathcal{T}$, and thus $\rho\in U$. Given $\sigma\in P(\rho)\cap N(\mathbb{R}^{3})$, we know $\rho+t\sigma\in R(\mathbb{R}^{3})\subset Y$. Since $Y$ is a topological vector space, we know when $t$ is small enough, $\rho+t\sigma\in U$. Therefore, $\rho+t\sigma\in R(\mathbb{R}^{3})\cap U=U_{R}$, and then $E_{J}(\rho)\leq E_{J}(\rho+t\sigma)$ for $t$ small enough. Then we know $\rho$ is a weak local minimizer restricted on $R(\mathbb{R}^{3})$. ∎
@@
Note that, unlike in McCann’s statement, here we first assume the local minimal energy $E_{J}(\rho)$ is finite, so that we can mimic the arguments in Lemma 5.3 and then make contradiction.
@@
Observing that A is a linear cone, it is intuitively reasonable to suspect that the minimum of f over A A may not exist. Therefore, we instead consider the minimization of f f over B B.
@@
Inspired by McCann’s Remark [31, Remark 3.7], we can also understand physically Proposition 5.15 as the following: suppose there is a small portion of mass in a rotating star which is moved to a very distant location, that is, from $\kappa=\rho(x)+\sigma(x)$ to $\kappa_{N}=\rho(x)+\sigma(x-N)$ for $N$ large. Then intuitively as $N\rightarrow\infty$, the internal energy $U$ and gravitational interaction energy $G$ will not change to much, while $I(\kappa_{N})\rightarrow\infty$, which implies the kinetic energy $T_{J}\rightarrow 0$, leading to the decrease of $E_{J}$ in the end.
@@
In McCann’s discussion [31, Remark 3.7], the conclusion considers $\rho$ as a local minimizer restricted to $R_{0}(\mathbb{R}^{3})$ rather than $R(\mathbb{R}^{3})$. However, as revealed by the proof above, if we focus solely on the restriction to $R_{0}(\mathbb{R}^{3})$, it becomes challenging to demonstrate the existence of sufficiently many functions $\sigma$ to ensure that the (EL) holds. Consequently, it is difficult to establish the non-existence of such $\rho$.
Remark 5.20.
From the proof above we notice that the issue comes from the fact that even if $\sigma(x)$ remains positive when $x$ is far away from spt $\rho$, $\rho+t\tilde{\sigma}$ is still continuous with respect to $t$, which implies local minimizer $\rho$ is also a weak local minimizer. Fortunately, this situation does not occur if ${R}\left(\mathbb{R}^{3}\right)$ is topologized via the Wasserstein $L^{\infty}$ metric. This is because if $\sigma(x)$ remains positive when $x$ is far away from spt $\rho$, then $\rho+t\tilde{\sigma}$ might not be continuous with respect to t, or $\rho+t\tilde{\sigma}$ might not fall within the neighborhood of $\rho$. In another word, a local minimizer is not necessary to be a weak local minimizer as we have seen in Remark 5.10.
From the proof of Proposition 5.15, we obtain the nonexistence of local minima with finite energy $E_{J}(\rho)$. The finiteness assumption is essential for following the arguments in Lemma 5.3 to compute derivative of $U(\rho)$. However, if the minimal energy $E_{J}(\rho)$ blows up, i.e., becomes infinite, the arguments in Lemma 5.3 can no longer be utilized.
One might hope to adopt an argument similar to Lemma 5.7 to show that if $\rho$ is a local minimum, its energy must necessarily be finite. While this is achievable under the topology induced by the Wasserstein $L^{\infty}$ distance, it remains uncertain whether such a $\rho$ with finite energy can be found under the topology inherited from the topological vector space. The following theorem leads to a conjecture that the statement regarding local minimizer (w.r.t. topology inherited from the topological vector space) having finite energy might not hold.
@@
If $P(\rho)$ in addition satisfies (F5), and $\rho\in R_{0}(\mathbb{R}^{3})$ satisfies $E_{J}(\rho)=\infty$. Then $\forall\sigma\in P(\rho)=\{\sigma\in X\mid\exists\epsilon(\sigma),\text{such that~}\rho+t\sigma\in\widetilde{U},0\leq t<\epsilon(\sigma)\}$, $\exists\delta>0$, such that $\forall t\in(0,\delta)$, we have $E_{J}(\rho+t\sigma)=\infty$. In particular, $\rho$ is a weak local minimizer (see Appendix A.1) for $E_{J}(\rho)$ on $R(\mathbb{R}^{3})$.
@@
If $E_{J}(\rho)=\infty$, due to Remark 5.1, we know $U(\rho)={\int{A(\rho)}}=\infty$. Since $\rho$ has compact support, $A(s)$ is an increasing function of $s$, and $\rho=0$ if and only if $A(\rho)=0$, then we know $\forall N>0$,
${\int_{\{{\rho\leq N}\}}{A(\rho)}}\leq A(N){\int_{\{{\rho\leq N}\}}1}\leq A(N){\int_{spt~\rho}1}=A(N)\mu\left({spt~\rho}\right)<\infty$
@@
Since we know ${\int{A(\rho)}}={\int_{\{{\rho\leq N}\}}{A(\rho)}}+{\int_{\{{\rho>N}\}}{A(\rho)}}=\infty$, we know $\int_{\{{\rho>N}\}}{A(\rho)}$$\infty$ for all $N>0$.
On the other hand, suppose $\sigma\in P(\rho)$, then we know $\forall t\in(0,\epsilon(\sigma))$, we have $\rho+t\sigma\geq 0$, then $\sigma\geq-\frac{\rho}{t}\geq-\frac{\rho}{\epsilon(\sigma)}$. If $t\rightarrow 0^{+}$, notice $A(s)$ is a increasing function, we have $U\left({\rho+t\sigma}\right)={\int{A(\rho+t\sigma)}}\geq{\int{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}}\geq{\int_{\{{\rho>N}\}}{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}}$ for all $N$. Since $t>0$, then $\lambda=1-\frac{t}{\epsilon(\sigma)}<1$, and ${\lim\limits_{t\rightarrow 0^{+}}\lambda}=1$.

### v2

@@ 5.2 Locally constant Lagrange multiplier
In this subsection, we essentially follow McCann’s arguments [31, Section 5] but with some refinements to prove “weak” versions of Theorem 2.22 Parts (iviv-$v$). The “weak” version serves as the bridge linking the norm of $\rho$ and the norm of the potential $V_{\rho}$, which can help to show the global continuity of the $W^{\infty}$ local minimizer $\rho$ as we have seen in the proof of Theorem 2.22
@@
We refer to [31], but slightly modify the arguments and provide a more detailed account (see Remark 5.8). Define the convex cone $P_{loc}:=\left\{\sigma\in P_{\infty}\mid\text{spt }\sigma\subset M\right\}$, which is a cone, and let $U=\{\rho\in X\mid\rho\geq 0,U(\rho)<\left.\infty,G(\rho,\rho)<\infty,T_{J}(\rho)<\infty\right\}$ be the same as in subsection 5.1. Thanks to Lemma 5.3 and Lemma 5.6, we have $\rho\in U,E_{J}(\rho)$ is $P_{loc}$-differentiable at $\rho$. Given $\kappa\in W_{loc}={R}\left(\mathbb{R}^{3}\right)\cap\{\rho+\sigma\in U\mid\sigma\in P_{loc}\}$, where the mass constraint is satisfied, thanks to Lemma 4.5 ($i$) we have $W^{\infty}(\rho,\kappa)\leq\operatorname{diam}M\leq 2\delta$, which shows that $\rho$ minimizes $E_{J}(\kappa)$ on $W_{loc}$. In particular, $\rho$ is a weak local minimizer restricted on $W_{loc}$ (see Appendix A for the definition). Moreover, since the open set $M$ intersects spt $\rho$, it must carry positive mass under $\rho$, see for example [17, Section 3.4, Exercise 25]. Therefore, similar to Remark 4.8, there is a smaller subset $C\subset M$ of positive measure on which $\rho(x)$ is bounded away from zero and infinity. Let $\mathbf{1}_{C}$ be the indicator function of set $C$, then by construction we have both $\pm\mathbf{1}_{C}\in P_{loc}$. For all $\sigma\in P_{loc}$, let $w=\sigma-\frac{\int_{\mathbb{R}^{3}}\sigma\,dx}{\int_{\mathbb{R}^{3}}\mathbf{1}_{C}\,dx}\mathbf{1}_{C}$, thanks to Remark 4.9, we know $\rho+tw\in W_{loc}$ when $t$ is small enough. Thanks to Theorem A.4 and Remark A.5, the conditions above imply that there is a unique Lagrange multiplier $\lambda\in R$ such that $\forall\sigma\in P_{loc}$, we have
$\int_{\mathbb{R}^{3}}E_{J}^{\prime}(\rho)\sigma\,dx\geq\lambda\int_{\mathbb{R}^{3}}\sigma\,dx$ (32)
Claim:
  1. (1)
    $E_{J}^{\prime}(\rho)(x)=A^{\prime}(\rho(x))-V_{\rho}(x)-\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))\geq\lambda\quad\text{on }M\text{ a.e.}$
  2. (2)
    $E_{J}^{\prime}(\rho)=\lambda\quad\text{on }M\cap\{x\mid\rho(x)>0\}\text{ a.e.}$
@@
In fact, if $E_{J}^{\prime}(\rho)<\lambda$ on a subset $K\subset M$ which has positive measure, similar to Remark 4.7, this subset may be taken slightly smaller so that $\rho$ is bounded on $K$. Then we take $\sigma=\mathbf{1}_{K}\in P_{loc}$ and would see it contradict (32), thus Claim (1) holds true. On the other hand, if $E_{J}^{\prime}(\rho)>\lambda$ on a subset $K\subset M$ with positive measure and $\rho>0$ in $K$, then similarly $K$ may be taken slightly smaller so that $\rho$ is bounded away from zero and infinity on $K$; in this case $\mathbf{1}_{K}\in P_{loc}$ can make a contradiction to (32) again, thus Claim (2) holds true.
@@
By (4) we know $A^{\prime}(0)=0$, and $A^{\prime}(s)>0$ when $s>0$. If $V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda>0$, then by Claim (1) $A^{\prime}(\rho(x))\geq V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda>0$, which implies $\rho>0$. By Claim (2) we can replace inequality by equation and get $A^{\prime}(\rho(x))=V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda$. If $V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\leq 0$, then $\rho=0$, otherwise $\rho>0$ implies again $0<A^{\prime}(\rho(x))=V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\leq 0$, which leads to a contradiction. Therefore, we have (EL), that is $A^{\prime}(\rho(x))=\left[V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\right]_{+}$ holds for almost all $x\in M$. Since $M$ has positive measure and $\rho$ in $M$ has positive mass, we know there exists a $x_{0}$ such that $\rho(x_{0})>0$ and thus $A^{\prime}(\rho(x_{0}))>0$, and (EL) holds at $x_{0}$. In particular, $0<A^{\prime}(\rho(x_{0}))=V_{\rho}(x_{0})+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x_{0}-\bar{x}(\rho))+\lambda$. Hence, it is clear that $\lambda$ is unique and given by this expression. ∎
@@
Compared to McCann’s proof, we have added the following discussions:
@@
Recall that in Section 4, we know $\rho+tw\in{R}\left(\mathbb{R}^{3}\right)$ may not be $W^{\infty}$-continuous as function of $t$. In other words, whether a local minimizer is also a weak local minimizer remains unclear in general. However, at least we know a minimizer restricted to $W_{loc}$ is a weak local minimizer restricted to $W_{loc}$, so that we can apply Theorem A.4 and Remark A.5. This is why we first consider the local case Proposition 5.7 (the diameter of $M$ is small) before addressing the next case Proposition 5.10.
@@
Given $y\in\Omega_{i}$, we know the ball $B_{\delta}(y)$ intersects spt $\rho$, where $B_{\delta}(y)$ is the open ball defined by $B_{\delta}(y):=\left\{x\in\mathbb{R}^{3}\mid|x-y|<\delta\right\}$. Thus Proposition 5.7 guarantees a unique $\lambda(y)$ such that (EL) holds a.e. on $B_{\delta}(y)$. The claim is that $\lambda(y)$ is independent of $y$. In fact, fix $y_{0}\in\Omega_{i}$, since $B_{\delta}(y_{0})$ is open, it will also be true that a slightly smaller ball $B_{\delta-\epsilon}(y_{0})$ intersects spt $\rho$. If $|x-y_{0}|<\epsilon$, then $M=B_{\delta}(x)\cap B_{\delta}(y_{0})$ intersects spt $\rho$ since $B_{\delta-\epsilon}(y_{0})\subset M$, and then we know $x\in\Omega_{i}$. $M$ has positive measure, thanks to Proposition 5.7, the uniqueness of $\lambda$ corresponding to $M$ forces $\lambda(x)=\lambda(y_{0})$. Thus $\lambda(y)$ is locally constant. As a result, the disjoint sets $C=\left\{x\in\Omega_{i}\mid\lambda(x)=\lambda(y_{0})\right\}$ and $D=\left\{x\in\Omega_{i}\mid\lambda(x)\neq\lambda(y_{0})\right\}$ are both open. Since $\Omega_{i}=C\cup D$ is connected and $y_{0}\in C$ implies $C$ is not empty, $C=\Omega_{i}$. Defining $\lambda_{i}:=\lambda(y_{0})$,(EL) must be satisfied a.e. on $\Omega_{i}$.
@@
Notice we discuss above the connected components of the $\delta$-neighbourhood of spt $\rho$. But the result can be easily converted to the case about the $\delta$-neighbourhood of connected components of spt $\rho$.
@@
From the above proof, we can see the power of Wasserstein metric, which can extend the Euler-Lagrange equation (EL) to a $\delta$-neighborhood of connected component of spt $\rho$, which help us to show the global continuity of $\rho$ and then (EP’) holds as we discuss in Theorem 2.22. Moreover, it plays a role in the process of proving the Lagrange multiplier is negative. Actually, there is another way to show Lagrange multipliers should be negative, given by Auchmuty and Beals [5], with different conditions.
@@ 5.3 Minimizers’ existence and nonexistence
In Section 4, we know $\rho+tw\in{R}\left(\mathbb{R}^{3}\right)$ may not be $W^{\infty}$-continuous as a function of $t$, and whether a local minimizer is also a weak local minimizer remains unclear in general. For comparison, we prove a local minimizer is a weak local minimizer under the topology inherited from a topological vector space. Furthermore, we demonstrate that, in this case, a local minimizer with finite energy does not exist.
Denote $X=L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, $\widetilde{U}=\{\rho\in X\mid\rho\geq 0\}$. We first show in the topology inherited from the topological vector space, a local minimizer is a weak local minimizer.
@@
If the topology on ${R}\left(\mathbb{R}^{3}\right)$ is inherited from a topological vector space, and $\rho$ is a local minimizer for $E_{J}(\rho)$ restricted on ${R}\left(\mathbb{R}^{3}\right)$, then $\rho$ is a weak local minimizer restricted on ${R}\left(\mathbb{R}^{3}\right)$, which is defined in Appendix A.
@@
We denote $\left.N\left(\mathbb{R}^{3}\right)=\left\{{\sigma\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)}\right|\int_{\mathbb{R}^{3}}\sigma\,dx=0\right\}$. We recall the setup in Appendix A, that $P(\rho)=\left\{{\sigma\in X}\right|\exists\epsilon(\sigma),~\text{such that}~\rho+t\sigma\in\widetilde{U},~0\leq t<\epsilon(\sigma)\}$.
Since the topology on $R(\mathbb{R}^{3})$ is inherited from a topological vector space, we know there is a topological vector space $Y$ with topology $\mathcal{T}$, such that $R(\mathbb{R}^{3})$ is a subset of $Y$, with the collection $\mathcal{T}_{R}=\left\{R\left(\mathbb{R}^{3}\right)\cap O\middle|O\in\mathcal{T}\right\}$ being the topology on $R(\mathbb{R}^{3})$. (See [33, Chapter 1, Section 16]).
Since $\rho$ is a local minimizer restricted on $R(\mathbb{R}^{3})$, then $\exists O_{R}\in\mathcal{T}_{R}$, such that $\rho\in O_{R}$ and $E_{J}(\rho)\leq E_{J}(\kappa)$ for all $\kappa\in O_{R}$. Notice that $O_{R}=R(\mathbb{R}^{3})\cap O$ for some $O\in\mathcal{T}$, and thus $\rho\in O$. Given $\sigma\in P(\rho)\cap N(\mathbb{R}^{3})$, we know $\rho+t\sigma\in R(\mathbb{R}^{3})\subset Y$. Since $Y$ is a topological vector space, we know when $t$ is small enough, $\rho+t\sigma\in O$. Therefore, $\rho+t\sigma\in R(\mathbb{R}^{3})\cap O=O_{R}$, and then $E_{J}(\rho)\leq E_{J}(\rho+t\sigma)$ for $t$ small enough. Then we know $\rho$ is a weak local minimizer restricted on $R(\mathbb{R}^{3})$. ∎
@@
Note that, unlike in McCann’s statement [31, Remark 3.7], here we first assume the local minimal energy $E_{J}(\rho)$ is finite, so that we can mimic the arguments in Lemma 5.3 and then make a contradiction.
@@
Inspired by McCann’s Remark [31, Remark 3.7], we can also understand physically Proposition 5.14 as the following: suppose there is a small portion of mass in a rotating star which is moved to a very distant location, that is, from $\kappa=\rho(x)+\sigma(x)$ to $\kappa_{N}=\rho(x)+\sigma(x-N)$ for $N$ large. Then, intuitively, as $N\rightarrow\infty$, the internal energy $U$ and gravitational interaction energy $G$ will not change too much, while $I(\kappa_{N})\rightarrow\infty$, which implies the kinetic energy $T_{J}\rightarrow 0$, leading to the decrease of $E_{J}$ in the end.
@@
In McCann’s discussion [31, Remark 3.7], the conclusion considers $\rho$ as a local minimizer restricted to $R_{0}(\mathbb{R}^{3})$ rather than $R(\mathbb{R}^{3})$. However, as revealed by the proof of Proposition 5.14, if we focus solely on the restriction to $R_{0}(\mathbb{R}^{3})$, it becomes challenging to demonstrate the existence of sufficiently many functions $\sigma$ to ensure that the (EL) holds. Consequently, it is difficult to establish the non-existence of such $\rho$.
@@
From the proof above we notice that the issue comes from the fact that even if $\sigma(x)$ remains positive when $x$ is far away from spt $\rho$, $\rho+t\tilde{\sigma}$ is still continuous with respect to $t$, which implies local minimizer $\rho$ is also a weak local minimizer. Fortunately, this situation does not occur if ${R}\left(\mathbb{R}^{3}\right)$ is topologized via the Wasserstein $L^{\infty}$ metric, because continuity of $\rho+t\tilde{\sigma}$ is not guaranteed under this topology, as discussed in Section 4.
From the proof of Proposition 5.14, we obtain the nonexistence of local minima with finite energy $E_{J}(\rho)$. The finiteness assumption is essential for following the arguments in Lemma 5.3 to compute the derivative of $U(\rho)$. However, if the minimal energy $E_{J}(\rho)$ blows up, i.e., becomes infinite, the arguments in Lemma 5.3 can no longer be utilized.
The following Proposition 5.21 leads to a conjecture that the statement regarding local minimizer (w.r.t. topology inherited from the topological vector space) having finite energy might not hold. Note that this proposition differs from Lemma 5.6, making a comparison interesting.
@@
If $P(\rho)$ in addition satisfies (F5), and $\rho\in R_{0}(\mathbb{R}^{3})$ satisfies $E_{J}(\rho)=\infty$. Then $\forall\sigma\in P(\rho)=\{\sigma\in X\mid\exists\epsilon(\sigma),\text{such that~}\rho+t\sigma\in\widetilde{U},0\leq t<\epsilon(\sigma)\}$, $\exists\delta>0$, such that $\forall t\in(0,\delta)$, we have $E_{J}(\rho+t\sigma)=\infty$. In particular, $\rho$ is a weak local minimizer (see Appendix A) for $E_{J}(\rho)$ on $R(\mathbb{R}^{3})$.
@@
If $E_{J}(\rho)=\infty$, due to Remark 5.1, we know $U(\rho)={\int_{\mathbb{R}^{3}}{A(\rho)}\,dx}=\infty$. Since $\rho$ has compact support, $A(s)$ is an increasing function of $s$, and $\rho=0$ if and only if $A(\rho)=0$, then we know $\forall N>0$,
${\int_{\{{\rho\leq N}\}}{A(\rho)}\,dx}\leq A(N){\int_{\{{\rho\leq N}\}}1\,dx}\leq A(N){\int_{spt~\rho}1\,dx}=A(N)\mu\left({spt~\rho}\right)<\infty$
@@
Since we know ${\int_{\mathbb{R}^{3}}{A(\rho)}\,dx}={\int_{\{{\rho\leq N}\}}{A(\rho)}\,dx}+{\int_{\{{\rho>N}\}}{A(\rho)}\,dx}=\infty$, we know $\int_{\{{\rho>N}\}}{A(\rho)}\,dx$$\infty$ for all $N>0$.
On the other hand, suppose $\sigma\in P(\rho)$, then we know $\forall t\in(0,\epsilon(\sigma))$, we have $\rho+t\sigma\geq 0$, then $\sigma\geq-\frac{\rho}{t}\geq-\frac{\rho}{\epsilon(\sigma)}$. If $t\rightarrow 0^{+}$, notice $A(s)$ is a increasing function, we have for all $N$,
$U\left({\rho+t\sigma}\right)={\int_{\mathbb{R}^{3}}{A(\rho+t\sigma)}\,dx}\geq{\int_{\mathbb{R}^{3}}{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}\,dx}\geq{\int_{\{{\rho>N}\}}{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}\,dx}$
Since $t>0$, then $\lambda=1-\frac{t}{\epsilon(\sigma)}<1$, and ${\lim\limits_{t\rightarrow 0^{+}}\lambda}=1$.